Prove that the map T: D to R^2 defined by T(x,y)= (u,v) is not invertible where u=x cos y , v= x sin y.
1
Expert's answer
2021-07-11T16:30:11-0400

T(x, y)= (u, v)


where u = x cos y and v = x sin y


To check whether the mapping T: D "\rightarrow" R2 is invertible or not we will find the Jacobian.



"J_T(x, y)" = "\begin{vmatrix}
 \dfrac{ ∂u}{ ∂x} & \dfrac{ ∂u}{ ∂y} \\ 
 \dfrac{ ∂v}{ ∂x} & \dfrac{ ∂v}{ ∂y}
\end{vmatrix}"



Now


"\dfrac{ ∂u}{ ∂x}" = cos y



"\dfrac{ ∂u}{ ∂y}" = - x sin y



"\dfrac{ ∂v}{ ∂x}" = sin y



"\dfrac{ ∂v}{ ∂y}" = x cos y




"J_T(x, y)" = "\begin{vmatrix}
 cos \ y & - x \ sin \ y \\ 
 sin \ y & x \ cos \ y
\end{vmatrix}"


"J_T(x, y)" = x cos2 y + x sin2 y



"J_T(x, y)" = x


Now the inverse of T(x, y)= (u, v)  will be the inverse of matrix "J_T(x, y)".  But according to the domain of T(x, y) x can be zero. So the function will not be invertible as matrix JT (x, y) will become singular . Hence, we say that T(x, y) = (u, v) is not invertible.